OpenStudy (pulsified333):

Let f and g be functions that satisfy f′(2)=−2 and g′(2)=11. Find h′(2) for each function h given below: (A) h(x)=6f(x). h′(2) = (B) h(x)=−5g(x). h′(2) = (C) h(x)=10f(x)+9g(x). h′(2) = (D) h(x)=8g(x)−11f(x). h′(2) = (E) h(x)=2f(x)+11g(x)−12. h′(2) = (F) h(x)=−2g(x)−3f(x)+10x. h′(2) =

OpenStudy (pulsified333):

@Michele_Laino

OpenStudy (michele_laino):

hint: case a) \[h'\left( x \right) = 6f'\left( x \right)\] case b) \[h'\left( x \right) = - 5g'\left( x \right)\] case c) \[h'\left( x \right) = 10f'\left( x \right) + 9g'\left( x \right)\] case d) \[h'\left( x \right) = - 11f'\left( x \right) + 8g'\left( x \right)\] case e) \[h'\left( x \right) = 2f'\left( x \right) + 11g'\left( x \right)\] case f) \[h'\left( x \right) = - 3f'\left( x \right) - 2g'\left( x \right) + 10\]